OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (pawanyadav):

Don't you think it's follows a pattern Every successive term is 9 more than previous term

OpenStudy (pawanyadav):

So it's an AP series

OpenStudy (pawanyadav):

Can you find total number of terms in this series.

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

wait is it 15?

OpenStudy (pawanyadav):

You know the formula to find last term of AP series

OpenStudy (anonymous):

no i dont

OpenStudy (pawanyadav):

The formula is Last term= first term + (n-1)d Where n=no. Of term and d=difference between two successive term

OpenStudy (anonymous):

is n 15

OpenStudy (pawanyadav):

Let's check by formula 135=-4 +(n-1)9 can you understand this

OpenStudy (anonymous):

wait why 135 isnt it 131

OpenStudy (pawanyadav):

Just respective values are plugged

OpenStudy (pawanyadav):

Yes right

OpenStudy (anonymous):

131=-4 +(n-1)9

OpenStudy (pawanyadav):

Can you solve it for n

OpenStudy (anonymous):

131=-4+9n-9 131=9n-13 9n=144 n=16

OpenStudy (pawanyadav):

Right

OpenStudy (pawanyadav):

Now you know the summation of AP series?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

summation of the (-4+9n) from n equals zero to fifteen

OpenStudy (pawanyadav):

Sorry that is (n/2)[2a+(n-1)d] OR (n/2)sum of first and last term

OpenStudy (pawanyadav):

You can do now

OpenStudy (anonymous):

what?

OpenStudy (pawanyadav):

Sum =16/2(-4+131) =8(128)= 1024